OpenStudy (anonymous):

please help i dont understand What would be the resulting molarity of a solution made by dissolving 31.3 grams of Ca(OH)2 in enough water to make a 1050-milliliter solution? Show all of the work needed to solve this problem.

OpenStudy (aaronq):

apply the molarity formula: \(M=\dfrac{n_{Solute}}{L_{Solution}}\)

OpenStudy (somy):

using gram given find mole of Ca(OH)2 mole= mass/ Mr then use formula Molarity= mole/dm3 so convert milliliters to dm3 and then substitute your readings to the equation that't it :)